OpenStudy (anonymous):

I am asking a question about "Complex Number" how this equation comes up: i^2 = (square root of -1)^2 = square root of (-1)^2 = square root of 1 = 1 ^ represents to the power. Anyone please answer.

OpenStudy (amistre64):

i^2 = -1 i = sqrt(-1)

OpenStudy (anonymous):

i = sqrt(-1)

OpenStudy (amistre64):

i = sqrt(-1) i^2 = [sqrt(-1)]^2 i^2 = -1

OpenStudy (amistre64):

[sqrt(-1)]^2 = -1^(2/2) = -1^1 = -1

OpenStudy (anonymous):

\[i ^{2}=1\] how is this possible?

OpenStudy (amistre64):

\begin{array}c i&i^2&i^3&i^4\\ i&-1&-i&1 \end{array} and the powers of I just keep repeating thru

OpenStudy (amistre64):

i^2 doesnt = 1

OpenStudy (amistre64):

your downfall is in trying to define i as a variable; when i is already defined

OpenStudy (anonymous):

go to this url : http://www.purplemath.com/modules/complex.htm here the equation : \[i ^{2}=\left( \sqrt{-1} \right)^{2}=\sqrt{\left( -1^{} \right)^{2}}=\sqrt{1}=1\] is given. how is this coming???

OpenStudy (amistre64):

they are not using "i" is a complex number; they are using it as a variable. i^2 does not equal sqrt[(-1)^2]

OpenStudy (anonymous):

i^2 isint 1...

OpenStudy (amistre64):

to wit they expalin in the next part: But this doesn't make any sense! You already have two numbers that square to 1; namely –1 and +1. And i already squares to –1. So it's not reasonable that i would also square to 1. This points out an important detail: When dealing with imaginaries, you gain something (the ability to deal with negatives inside square roots), but you also lose something (some of the flexibility and convenient rules you used to have when dealing with square roots). In particular, YOU MUST ALWAYS DO THE i-PART FIRST!

OpenStudy (anonymous):

lol that was a very complicated way of saying that

OpenStudy (amistre64):

:) thats the purple math for you :) they give you a proof by absurdity

OpenStudy (anonymous):

@amistre thank you.

OpenStudy (anonymous):

hold up, i realized something though, check this out:

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

bye everyone

OpenStudy (anonymous):

\[i ^{2} =(\sqrt{-1})^{2}= (\sqrt{-1})(\sqrt{-1}) = \sqrt{-1*(\sqrt{-1})^{2}}=\sqrt{-1*(repeat)}\] at which point you would have to define that either \[(\sqrt{a})^{2}=a \] or \[(\sqrt{a})^{2}=\sqrt{a ^{2}}=\pm a\]

OpenStudy (anonymous):

im guessing for imaginaries they defined the former rather than the latter

OpenStudy (radar):

Do not agree with the 4th expression of the first line.\[(\sqrt{-1})(\sqrt{-1})\neq \sqrt{-1*(\sqrt{-1)^{2}}}\]